# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v6

@@ III-B Visual Attention Depth Net
III-C Auto-learning uncertainty temporal-spatial fusion
When testing our designed VADepth model, we found that the Auto-masking[9] method almost failed, as shown in Fig.7. This issue is more severe in the Cityscapes[11]. Upon inspecting the output of Auto-masking, we discovered that it couldn’t perfectly shield non-rigid objects. We attribute this to the model’s robustness, the limitations of the Auto-masking method, and the complexity of the Cityscapes[11, 55].
Therefore, we need to address the issue of non-rigid objects to further enhance performance. In temporal sequential frames, there are freely moving targets, such as pedestrians and vehicles(as show in Fig.8). These do not conform to the rigid scene transformation assumptions based on depth and camera pose. The solution provided by Auto-Masking[9] is to compare the pixels before and after reprojection. If the reprojection error after reprojection is worse than without reprojection, it indicates that the pixel belongs to a non-rigid object.
$\mu=[\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})<\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c})].$ (8)
However, the network can still rationalize the depth of non-rigid moving objects by pulling forward targets further away and pushing backward closer targets, making them consistent with the camera extrinsics(Pose), as show in Fig.7. In fact, non-rigid moving objects only exist in temporally related frames, while in spatially related frames (stereo settings), since they are captured at the same moment, this issue does not arise. Therefore, introducing stereo supervision into the system is the most direct and effective method, rather than merely excluding masks for non-rigid objects[55, 9].
Figure 9: Auto-Spatio-Temporal Uncertainty Fusion: The Uncertainty Generation Network utilizes a Pose Encoder and establishes an Uncertainty Decoder. It inputs the $\textbf{I}_{\{-1,+1,0,s\}}$ and the target frame $\textbf{I}_{0}$ separately to output the uncertainty map $\hat{\textbf{U}}$. Then, it generates the weight map U for different error maps through a “1-Softmax” operation.
But when comparing temporal sequence reprojection with spatial reprojection, each has its own advantages and disadvantages. First, the translation of pose between temporally related frames are larger, whereas in spatial reprojection (stereo), the baseline is fixed and small. This makes temporal-based transformations better suited for estimating distant targets. Then, temporally related frames typically include three frames: the current frame and the preceding and following frames, allowing information from the preceding and following frames to fill in occluded areas in the target frame. In spatial reprojection (stereo), regardless of whether the target frame is set as the left or right eye, there is no symmetric completion of occluded areas. Thus, we seek a method that can effectively combine the advantages of both temporal and spatial reprojections to predict better depth maps.
We referenced prior studies on Uncertainty Learning[56, 57, 22]. And in the work by Zhu et al.[22], a method was introduced to address occlusion problems by automatically predicting uncertainty and reweighting the reprojection in the reprojection process of spatially adjacent frames. Building on this idea, we aim to take it a step further by introducing a method to predict uncertainty during both temporal and spatial reprojections. This approach seeks to leverage the strengths of both temporal and spatial domains effectively.
To combine photometric loss across different frames, Zhu et al.[22] specify $\sum_{c}(1-\hat{u}_{i,c})=1,c\in\{t-1,t+1\}$. Hence, they combine the output uncertainty map $\hat{\textbf{U}}_{c}(\hat{\textbf{U}}_{c}\geq 0)$ with softmax: $\{\textbf{U}_{t-1},\textbf{U}_{t+1}\}=1-Softmax\{\hat{\textbf{U}}_{t-1},\hat{\textbf{U}}_{t+1}\}$, to ensure that $\sum_{c}u_{i,c}=1$. The final joint optimization function is written as:
$\begin{split}\mathcal{L}&=\frac{1}{N}\sum^{N}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})\\ &\quad+\alpha\frac{1}{N}\sum^{N}\sum_{c}\hat{u}_{i,c},\quad\sum_{c}u_{i,c}=1.\end{split}$ (9)
Using this formula, the network will automatically learn the uncertain regions(areas where the loss cannot be reduced through optimization) within the loss, forming the final weights for multi-frame loss through comparison. In the temporal multi-frame reprojection loss combination, high-uncertainty regions are the frames’ occluded areas revealed due to reprojection. Meanwhile, in the spatial and temporal reprojection loss combination, high-uncertainty regions in the temporal reprojection loss map are the non-rigid motion areas that can be accurately reprojected in the spatial.
$\begin{split}\mathcal{L}^{uD}=\frac{1}{N}\sum^{N}\Bigg{(}u_{i,\tau}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})+u_{i,s}pe_{i}(\textbf{I}_{t},\textbf{I}_{s\rightarrow t})\Bigg{)}\\ +\alpha\frac{1}{N}\sum^{N}(\sum_{c}\hat{u}_{i,c}+\sum_{\sigma}\hat{u}_{i,\sigma}),\quad c\in\{-1,+1\},\quad\sigma\in\{\tau,s\}\end{split}$ (10)
where $\tau$ and $s$ represent the temporal domain and the spatial domain, respectively, with $\alpha$ set to $1e^{-2}$. $u_{i,\tau}$ and $u_{i,s}$ represent the temporal-domain uncertainty and the spatial-domain uncertainty, $\sum_{c}u_{i,c}=1,\sum_{\sigma}u_{i,\sigma}=1$, $\{\textbf{U}_{\tau},\textbf{U}_{s}\}=1-Softmax\{\hat{\textbf{U}}_{\tau},\hat{\textbf{U}}_{s}\}$. $s$ represents the source frame in a stereo setup. Since the Kitti dataset randomly selects either the left or right camera image as the target frame, the other frame becomes the source frame.
@@ III-C Auto-learning uncertainty temporal-spatial fusion
Regarding how to obtain temporal and spatial uncertainty maps, we consider referencing the method used to generate temporal uncertainty maps by inputting related frames containing uncertainty information. For spatial reprojection, we can input $\textbf{I}_{s}$ and $\textbf{I}_{0}$, but there is no single related frame that uniquely represents temporal reprojection. We attempt to input $\textbf{I}_{0}$ and $\textbf{I}_{0}$ as representatives. By comparing the output uncertainty maps when calculating the loss, we can form reasonable uncertainty maps, as shown in Fig. 9. Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues. We improved this method to fuse temporal and spatial reprojection errors. Since non-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty(Fig.10). Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process. By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depths.
III-D Training Strategy
III-D1 Synthetic supervised depth loss
@@
III-D2 Optimization Strategy
Swami et al.[43] highlighted that multiple networks and optimization losses in unsupervised domain adaptation methods often compete against each other and fail to contribute effectively to the optimization of the depth estimation. This lack of cooperation among different components makes it challenging to achieve best results. So here, we have set two optimizers $\{\mathcal{O}_{d2c},\mathcal{O}_{d2c}\}$ to separately train the Depth Net and the Color net, as shown in Fig.2.
Optimizer 1-$\mathcal{O}_{c2d}$: Supervised training of the Depth Net on the Syn-Real samples and the Color Net on the Real world unsupervised depth dataset for predicting color from depth, $\gamma=1e^{-3}$.
$\mathcal{L}^{\mathcal{O}_{d2c}}=\mathcal{L}^{d2c}+\mathcal{L}^{sD}+\gamma\mathcal{L}^{sm}.$ (13)
Optimizer 2-$\mathcal{O}_{c2d}$: Unsupervised training of the Depth Net on the Real world and applyed the background textureless regularization loss , $\beta=1e^{-2},\gamma=1e^{-3}$.
$\mathcal{L}^{\mathcal{O}_{c2d}}=\mathcal{L}^{uD}+\beta\mathcal{L}^{sm}+\gamma\mathcal{L}^{sky}.$ (14)
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.I, Tab.II, and Tab.IV). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.VIII). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 13, Fig. 15, and Fig. 16).
@@ IV-A Datasets
We utilized two unsupervised monocular depth estimation datasets: KITTI[25] and Cityscapes[11], along with synthetic datasets vKitti2[12], SHIFT[23], and UrbanSyn[24]. KITTI contains 39,810 training triplets, 4,424 for validation, and 697 for evaluation, featuring diverse scenes with rich colors and occasional non-rigid objects. Cityscapes, designed for autonomous driving, provides 69,731 monocular triplets from the LeftImg8bit Sequence for training, mainly urban scenes with non-rigid pedestrians and vehicles, often in darker tones. The baseline for KITTI stereo data is 54 cm, while for Cityscapes it is 22 cm. vKitti2, an upgraded version of Virtual KITTI, uses the Unity game engine for a more photorealistic environment and provides 21,260 frames for training. SHIFT features various weather conditions and times, offers dense depth information with 1 mm resolution, and includes simulated pedestrians. We selected 153,000 frames and synthetic depths from its front-view cameras. UrbanSyn, designed to address the synth-to-real domain gap, provides 7,539 images with depth labels, offering a highly realistic synthetic dataset for various computer vision tasks.
To leverage the variety of multiple synthetic datasets, we normalized and shuffled the specifications (size, depth resolution, etc.) of the vKitti2, SHIFT, and UrbanSyn datasets to create a unified dataset called MixSyn. The MixSyn dataset contains significantly more samples compared to KITTI and Cityscapes. Therefore, during the training process, MixSyn is trained once, while KITTI or Cityscapes is trained for multiple epochs, as shown in Fig.12. We demonstrate the advantages of using MixSyn in Tab. VII.
@@
Method Sup. Size Error Metric Accuracy Metric
@@
Monodepth2[9] m 192x640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
CADepth[29] m 192x640 0.105 0.769 4.535 0.181 0.892 0.964 0.983
DIFFNet[31] m 192x640 0.102 0.749 4.445 0.179 0.897 0.965 0.983
MonoFormer[60] m 192x640 0.106 0.839 4.627 0.183 0.889 0.962 0.983
Lite-Mono[1] m 192x640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
MonoViT[2] m 192x640 0.099 0.708 4.372 0.175 0.900 0.967 0.984
Back2Color m 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
DIFFNet[31] m 320x1024 0.097 0.722 4.345 0.174 0.907 0.967 0.984
Lite-[1] m 320x1024 0.097 0.710 4.309 0.174 0.905 0.967 0.984
MonoViT[2] m 320x1024 0.096 0.714 4.292 0.172 0.908 0.968 0.984
Back2Color m 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
@@ IV-B Implementation Details
Our training is implemented by Pytorch[61] with NVIDIA RTX 3090Ti(24GB). We have used ADAM[62] as the optimizer. We have set the learning rate to $1\times 10^{-4}$, the same as in Monodepth2[9], and the learning rate will decay to $1\times 10^{-5}$ after 15 epochs. The other parameters are the same as $(\beta_{1}$, $\beta_{2})$ = $(0.9,0.999)$. We set the SSIM weight to $0.85$ and the smoothness weight to $\alpha=1e-3$. Furthermore, we utilized a pretrained segmentation network[52] trained on the Cityscapes semantic segmentation dataset. We evaluate the performance using depth estimation metrics[25, 63]: rmse(log), rmse, Sq.relative, Abs relative, and Accuracy.
@@
Method Sup. Size Error Metric Accuracy Metric
@@
Monodepth2[9] ms 192x640 0.106 0.818 4.750 0.196 0.874 0.957 0.979
HR-Depth[30] ms 192x640 0.107 0.785 4.612 0.185 0.887 0.962 0.982
CADepth[29] ms 192x640 0.102 0.752 4.504 0.181 0.894 0.964 0.983
DIFFNet[31] ms 192x640 0.101 0.749 4.445 0.179 0.898 0.965 0.983
MonoViT[2] ms 192x640 0.098 0.683 4.333 0.174 0.904 0.967 0.984
Back2Color ms 192x640 0.092 0.685 4.229 0.170 0.911 0.968 0.984
Monodepth2[9] ms 320x1024 0.106 0.806 4.630 0.193 0.876 0.958 0.980
HR-Depth[30] ms 320x1024 0.101 0.716 4.395 0.179 0.894 0.966 0.983
CADepth[29] ms 320x1024 0.096 0.694 4.264 0.173 0.908 0.968 0.984
DIFFNet[31] ms 320x1024 0.094 0.678 4.250 0.172 0.911 0.968 0.984
MonoViT[2] ms 320x1024 0.093 0.671 4.202 0.169 0.912 0.969 0.984
Back2Color ms 320x1024 0.087 0.644 4.079 0.166 0.920 0.970 0.985
@@
Method Sup. Error Metric Accuracy Metric
@@ IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.I and Tab.II with Kitti Eigen Split[25] and Tab.IV with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.13 and Fig.15. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[1, 2] with larger params and more complex structures(Tab.VIII).
Figure 14: Temporal-Spatial Uncertainty Fusion: White represents higher reliability in spatial reprojection, while black represents higher reliability in temporal reprojection, as shown in the legend in the bottom right corner. We can observe that non-rigid moving objects in the image tend to reproject more in space (white). Meanwhile, occluded areas and distant regions tend to reproject more in the temporal sequence (black), and the remaining areas show no significant tendency (gray).
TABLE IV: Experiments on the Cityscapes benchmark with 192$\times$640. Methods* requiring semantic information and Back2Color requiring stereo set.
@@
MonoDEVSNet[18] 192x640 0.102 0.685 4.303 0.178 0.894 0.966 0.984
Back2Color 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
Swami et al.[43] 320x1024 0.103 0.654 4.300 0.178 0.891 0.966 0.984
Back2Color 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985

### v7

@@ III-B Auto-learning Uncertainty Temporal–Spatial Fusion
Inspired by uncertainty learning [68, 69, 44], we introduce an Auto-learning Uncertainty Temporal–Spatial Fusion (Auto-UTSF) mechanism to combine temporal and spatial reprojection losses.
In this section, subscripts $t$, t1t\!-\!1, t+1t\!+\!1, and $s$ are used to index temporal and stereo frames, while dataset/domain subscripts (e.g., $\mathcal{X}$ and $\mathcal{S}$) are omitted for clarity when the context is unambiguous. And $i$ indexes individual pixels in the image, and $N$ denotes the total number of pixels. Following [69], pixel-wise temporal uncertainty logits $\hat{u}^{\tau}_{i,k}$ are predicted for temporally adjacent frames $k\in\{t-1,t+1\}$ and normalized using a softmax operation:
$u^{\tau}_{i,k}=1-\text{Softmax}(\hat{u}^{\tau}_{i,t-1},\hat{u}^{\tau}_{i,t+1}),\quad\sum_{k}u^{\tau}_{i,k}=1.$ (6)
And The temporal reprojection loss is then defined as
$\mathcal{L}_{\tau}=\frac{1}{N}\sum_{i}\sum_{k}u^{\tau}_{i,k}\,pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})+\alpha\frac{1}{N}\sum_{i}\sum_{k}\hat{u}^{\tau}_{i,k},$ (7)
where $\alpha$ is set to $1\times 10^{-2}$.
To further fuse temporal and spatial supervision, we introduce domain-level uncertainty to balance their contributions. Temporal-domain and spatial-domain uncertainty logits $(\hat{u}^{\tau}_{i},\hat{u}^{s}_{i})$ are normalized as
$u^{\tau}_{i},\;u^{s}_{i}=1-\text{Softmax}(\hat{u}^{\tau}_{i},\hat{u}^{s}_{i}),\quad u^{\tau}_{i}+u^{s}_{i}=1.$ (8)
The final uncertainty-aware temporal–spatial fusion loss is formulated as
$\mathcal{L}_{\text{ud}}=\frac{1}{N}\sum_{i}\Big($ $u^{\tau}_{i}\sum_{k}u^{\tau}_{i,k}\,pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})$ (9)
$+u^{s}_{i}\,pe(\mathbf{I}_{t},\mathbf{I}_{s\rightarrow t})\Big)$
$+\alpha\frac{1}{N}\sum_{i}\Big(\sum_{k}\hat{u}^{\tau}_{i,k}+\hat{u}^{\tau}_{i}+\hat{u}^{s}_{i}\Big),$
where $\mathbf{I}_{s}$ denotes the stereo source image corresponding to the target frame $\mathbf{I}_{t}$.
Figure 7: Auto-Spatio-Temporal Uncertainty Fusion: Given temporal and stereo frame pairs, the uncertainty generation module predicts pixel-wise uncertainty logits, which are normalized via a $1-\mathrm{Softmax}$ operation to adaptively weight temporal and spatial reprojection errors.
Through uncertainty learning, the model automatically assigns low weights to unreliable regions in each domain. In temporal reprojection, high uncertainty mainly corresponds to occlusions and non-rigid motion, while in spatial reprojection, uncertainty primarily arises from occluded regions caused by viewpoint changes. By fusing uncertainties from both temporal and spatial domains, Auto-UTSF selectively suppresses unreliable supervision and exploits complementary geometric cues, enabling robust depth estimation in dynamic traffic scenes.
@@
III-C Training Strategy
III-C1 Synthetic supervised depth loss
@@
III-C2 Optimization Strategy
Swami et al. [52] pointed out that, in unsupervised domain adaptation methods, multiple networks and optimization objectives often compete with each other, thereby failing to jointly contribute to effective depth estimation. Such a lack of cooperation among different components makes optimization challenging and limits overall performance.
To address this issue, we adopt two separate optimizers to decouple the training process and independently optimize the Depth Network on synthetic data and real-world unsupervised data, respectively, as illustrated in Fig. 2. Each training batch contains both real-world samples for unsupervised depth estimation and synthetic color samples.
Optimizer Step I: The Depth Network is trained in an unsupervised manner on real-world samples using our proposed Auto-UTSF mechanism, while the Color Network is trained in a supervised manner on real-world data to predict color from depth. We set $\gamma=10^{-3}$, and $\mathcal{L}^{\text{sm}}$ denotes the smoothness loss. The overall loss for this step is defined as
$\mathcal{L}_{\mathcal{O}_{1}}=\mathcal{L}_{\text{d2c}}+\mathcal{L}_{\text{ud}}+\gamma\mathcal{L}^{\text{sm}}.$ (12)
Optimizer Step II: The Depth Network is further trained in a supervised manner using the proposed Syn-Real CutMix samples. The corresponding optimization objective is given by
$\mathcal{L}_{\mathcal{O}_{2}}=\mathcal{L}_{\text{sd}}+\gamma\mathcal{L}^{\text{sm}}.$ (13)
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.I, Tab.II, and Tab.IV). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VIII, Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.IX). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 11, Fig. 12, and Fig. 13).
@@ IV-A Datasets
We utilized two unsupervised monocular depth estimation datasets: KITTI[10] and Cityscapes[7], along with synthetic datasets vKITTI2[4], SHIFT[51], and UrbanSyn[gómez2023one]. KITTI contains 39,810 training triplets, 4,424 for validation, and 697 for evaluation, featuring diverse scenes with rich colors and occasional non-rigid objects. Cityscapes, designed for autonomous driving, provides 69,731 monocular triplets from the LeftImg8bit Sequence for training, mainly urban scenes with non-rigid pedestrians and vehicles, often in darker tones. The baseline for KITTI stereo data is 54 cm, while for Cityscapes it is 22 cm. vKITTI2, an upgraded version of Virtual KITTI, uses the Unity game engine for a more photorealistic environment and provides 21,260 frames for training. SHIFT features various weather conditions and times, offers dense depth information with 1 mm resolution, and includes simulated pedestrians. We selected 153,000 frames and synthetic depths from its front-view cameras. UrbanSyn, designed to address the synth-to-real domain gap, provides 7,539 images with depth labels, offering a highly realistic synthetic dataset for various computer vision tasks.
To leverage the variety of multiple synthetic datasets, we normalized and shuffled the specifications (size, depth resolution, etc.) of the vKITTI2, SHIFT, and UrbanSyn datasets to create a unified dataset called MixSyn. The MixSyn dataset contains significantly more samples compared to KITTI and Cityscapes. Therefore, during the training process, MixSyn is trained once, while KITTI or Cityscapes is trained for multiple epochs, as shown in Fig.10. We demonstrate the advantages of using MixSyn in Tab. VIII.
@@
Method Sup. [^1] Sup.(Supervision) in the tables means what trainning set is used, Video(m),Stereo(s). Size Error Metric Accuracy Metric
@@
[…]
@@ IV-B Implementation Details
All experiments are implemented using PyTorch [45] and conducted on a single NVIDIA RTX 3090Ti GPU with 24 GB memory. We adopt the Adam optimizer [31] with an initial learning rate of $1\times 10^{-4}$, following the setting of Monodepth2 [14]. The learning rate is decayed to $1\times 10^{-5}$ after 15 epochs, while the remaining hyperparameters are set to $(\beta_{1},\beta_{2})=(0.9,0.999)$. The SSIM loss weight is set to $0.85$, and the smoothness regularization weight is set to $\alpha=1\times 10^{-3}$. Depth estimation performance is evaluated using standard metrics [10], including $\mathrm{AbsRel}$, $\mathrm{SqRel}$, $\mathrm{RMSE}$, $\mathrm{RMSE}_{\log}$, and accuracy thresholds.
To ensure computational efficiency and real-time applicability in urban driving scenarios, we adopt a lightweight CNN-based backbone for feature extraction, specifically the Visual Attention Network (VAN) [19], which captures long-range dependencies with low computational cost. A minor architectural adaptation is introduced to preserve higher-resolution features for dense depth prediction, while keeping the original backbone structure unchanged. The extracted multi-level features are subsequently processed by a high-resolution depth decoder [22] to recover fine-grained spatial details and generate depth maps at the input resolution. Experimental results demonstrate that the proposed framework consistently outperforms commonly used CNN-based backbones, such as ResNet-34, HRDepth, and DIFFNet, and achieves competitive performance compared to Transformer-based methods [63, 62] with significantly lower computational complexity. VAN0-VAN3 denote four variants of the VAN backbone with different parameter scales, as summarized in TableIX. Although VAN is adopted as the default backbone for efficiency, Experimental results demonstrate that the proposed framework consistently improves performance when integrated with commonly used CNN-based backbones, such as ResNet-34, HRDepth, and DIFFNet, and achieves competitive performance compared to Transformer-based methods [63, 62] with significantly lower computational complexity, as shown in Fig.IX.
@@
Method Sup.11footnotemark: 1 Size Error Metric Accuracy Metric
@@
[…]
@@
Method Sup.11footnotemark: 1 Error Metric Accuracy Metric
@@ IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.I and Tab.II with KITTI Eigen Split[10] and Tab.IV with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.11 and Fig.12. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[62, 63] with larger params and more complex structures(Tab.IX).
[…]
@@
[…]